# Learning in a Multifield Coherent Ising Machine

### v2

@@ Discussion
We have shown that a metamaterial composed of a network of parametrically-driven, multimodal resonators can learn from examples—autonomously acquiring the capability to solve a classification task. The metamaterial training is based on a contrastive learning rule, which involves comparing two copies of the system—one left free, and one clamped to the desired output—and updating the model weights according to the difference between copies. However, unlike traditional contrastive learning, our model weights are binary and cannot be updated continuously. We circumvent this limitation by reformulating the learning rule as a probabilistic process, where the probability of a weight flip is a function of the mismatch between the two copies. This mismatch-dependent flip rate is reminiscent of evolutionary mechanisms in bacteria, where the mutation rate increases when the organism is under stress [34, 35, 36]. Similarly, our system increases the weight mutation rate at sites where its clamped and free copies disagree.
@@
By providing a tractable physical model for emergent learning—analogous to the tight-binding Hamiltonians that describe topological insulators and superconductors—our work invites condensed matter-inspired questions about learning. For instance, one can now investigate the influence of disorder [43, 44], aging [45], and phase transitions [46, 47, 48] on learning behavior. As the constituent oscillators can be realized with components that operate at optical frequencies [49], consume little energy [50], or are integrated with very high densities [51], our model may inspire new directions to train machine learning models with physical computers.
@@ Appendix A Determination of the parametric resonance regions
In section Single-site learning dynamics we discussed how detuning of the parametrically pumped Ising mode leads to conditional forgetting. Here, we provide an in-depth analysis of parametric resonance using a combination of Floquet theory and perturbation methods. Our main objectives are to derive the resonance curve shown in Fig. 2a and to show that detuning has the effect of shifting it. The relevant equation is that of the Ising mode (Eq. 2) without interactions and noise, which can be cast in the following form:
@@
for which the resonance regions can be found using Floquet theory. They form tongues emanating from the points $\delta=n^{2}/4$ along the $\delta$-axis in what is known as the Ince-Strutt diagram [52]. For the purposes of this paper, we consider the resonance region around the so-called parametric resonance condition, which holds when the driving frequency is twice the mode’s natural frequency, or equivalently when $\delta=1/4$. We can use perturbation theory to find an approximate expression for this resonance curve in the case where $\mu,\gamma>0$, following an approach adapted from [52] known as the the method of multiple scales. Firstly, we treat the pumping amplitude $\epsilon$ as a small perturbation, and rescale $\mu$ and $\gamma$ to be of the same order, $\mu=:\epsilon\tilde{\mu}$ and $\gamma=:\epsilon\tilde{\gamma}$. Next, we expand the resonance curve in powers of $\epsilon$, $\delta=1/4+\delta_{1}\epsilon+\dots$ and define two independent timescales; a fast one $\xi:=t$ and a slow one $\eta:=\epsilon t$. Because the timescales are assumed to be independent, the full time derivative with respect to $t$ splits into partial derivatives with respect to $\xi$ and $\eta$ as follows:

### v3

@@ Discussion
We have shown that a metamaterial composed of a network of parametrically-driven, multimodal resonators can learn to solve a classification task, when excited with example signals. The material performs its own weight update according to a learning rule, requiring only two simple drive clock signals $\lambda(t)$ and $\mu(t)$ in addition to the data and labels. However, the metamaterial does not require independent controllers adjusting each individual weight. The metamaterial learning dynamics is based on contrastive learning, which involves comparing two copies of the system—one left free, and one clamped to the desired output—and updating the model weights according to the difference between copies. However, unlike traditional contrastive learning, our model weights are binary and cannot be updated continuously. We circumvent this limitation by reformulating the learning rule as a probabilistic process, where the probability of a weight flip is a function of the mismatch between the two copies. This mismatch-dependent flip rate is reminiscent of evolutionary mechanisms in bacteria, where the mutation rate increases when the organism is under stress [35, 36, 37]. Similarly, our system increases the weight mutation rate at sites where its clamped and free copies disagree.
@@
By providing a tractable physical model for emergent learning—analogous to the tight-binding Hamiltonians that describe topological insulators and superconductors—our work invites condensed matter-inspired questions about learning. For instance, one can now investigate the influence of disorder [44, 45], aging [46], and phase transitions [47, 48, 49] on learning behavior. Although the present work is concerned with non-dimensional tight-binding models and thus no specific performance claims can be made, our model may inspire new directions to train machine learning models with physical computers, as the constituent oscillators can be realized with components that operate at optical frequencies [50], consume little energy [51], or are integrated with very high densities [52]. Our work also poses novel exciting fundamental questions, for example involving learning dynamics in discrete spin systems, architectural and scaling considerations, as well as addressing the experimental realization of such models.
@@ Appendix A Determination of the parametric resonance regions
In the manuscript we discussed how detuning of the parametrically pumped Ising mode leads to conditional forgetting. Here, we provide an in-depth analysis of parametric resonance using a combination of Floquet theory and perturbation methods. Our main objectives are to derive the resonance curve shown in Fig. 2a and to show that detuning has the effect of shifting it. The relevant equation is that of the Ising mode (Eq. 2) without interactions and noise, which can be cast in the following form:
@@
for which the resonance regions can be found using Floquet theory. They form tongues emanating from the points $\delta=n^{2}/4$ along the $\delta$-axis in what is known as the Ince-Strutt diagram [53]. For the purposes of this paper, we consider the resonance region around the so-called parametric resonance condition, which holds when the driving frequency is twice the mode’s natural frequency, or equivalently when $\delta=1/4$. We can use perturbation theory to find an approximate expression for this resonance curve in the case where $\mu,\gamma>0$, following an approach adapted from [53] known as the method of multiple scales. Firstly, we treat the pumping amplitude $\epsilon$ as a small perturbation, and rescale $\mu$ and $\gamma$ to be of the same order, $\mu=:\epsilon\tilde{\mu}$ and $\gamma=:\epsilon\tilde{\gamma}$. Next, we expand the resonance curve in powers of $\epsilon$, $\delta=1/4+\delta_{1}\epsilon+\dots$ and define two independent timescales; a fast one $\xi:=t$ and a slow one $\eta:=\epsilon t$. Because the timescales are assumed to be independent, the full time derivative with respect to $t$ splits into partial derivatives with respect to $\xi$ and $\eta$ as follows:
@@ Appendix B Numerical methods
.1 Simulation of the clamped boundary conditions
To impose the boundary conditions, we divide the degrees of freedom into internal (subscript $I$) and output (subscript $o$) components. We write the system $\mathbf{D}\mathbf{u}=\mathbf{f}$ in block form as:
$\begin{bmatrix}\mathbf{D}_{II}&\mathbf{d}_{Io}\\ \mathbf{d}_{oI}^{T}&d_{oo}\end{bmatrix}\begin{bmatrix}\mathbf{u}_{I}\\ u_{o}\end{bmatrix}=\begin{bmatrix}\mathbf{f}_{I}\\ f_{o}\end{bmatrix}$ (13)
where $u_{o}$ is the prescribed output degree of freedom, $f_{o}$ is the constraint force, and $F_{I}$ is the external force (which inputs the data into the system). From the first block row, we bring the prescribed output to the right-hand side to obtain the reduced system $\mathbf{D}_{II}\mathbf{u}_{I}=\mathbf{f}_{I}-\mathbf{d}_{Io}u_{o}$, which we solve for the unknown internal DOFs $\mathbf{u}_{I}$. Note that in the reduced system equation, it is not necessary to explicitly calculate the constraint force $f_{o}$ to simulate the system dynamics.
In an experimental setting, the boundary conditions could be applied by connecting the output degree of freedom to a very large mass, or by applying external feedback.
Disorder and defect site tolerance
For our proposed learning mechanism to be accessible in a future experiment, it must be able to survive under the presence of realistic fabrication variability. To demonstrate that this is indeed the case, we simulated how the learning performance on the Iris task is affected by the presence of two different types of imperfections, frequency disorder and damaged resonators. For each type of imperfection, we conducted 16000 simulations, each of them with a different random perturbation (either a different frequency perturbation or a different choice of stuck resonators). We modelled the frequency disorder by adding a random, independent Gaussian increment to the frequency of each resonator (Fig. 5a), and simulated stuck resonators by prescribing their associated spin state to zero (Fig. 5b).
Figure 5: Sensitivity to disorder. (a) Classification accuracy on the Iris dataset as a function of the resonator frequency disorder. (b) Classification accuracy on the Iris dataset as a function of the fraction of defective sites. Both results have been obtained by averaging 16000 simulations. The shaded area corresponds to one standard deviation.
Simulations reveal that learning is not appreciably impaired by disorders below 250 ppm (parts-per-million) or by a fraction of defective resonators below 2.5%. These numbers fall within the accessible experimental range: Resonators with frequency disorders of a few tens of ppm or better are routinely attainable in photonics [55] and mechanics [56]; similarly, defect free arrays of 50-100 devices are also routinely produced in both domains [57, 58]. These results provide strong evidence that, while the experimental realization of self-learning oscillator networks represents a significant experimental endeavor, it does not rely on unrealistic fabrication tolerances.
Operation under a square learning protocol
Throughout the paper, we computed the transition probabilities assuming the learning pulse $\mu(t)$ followed a Gaussian function. Generating such pulses may require additional circuitry, therefore, it is worth examining whether a simpler function, such as a square pulse, can achieve the same results. Figure 6 compares the transition probabilities obtained by switching the learning potential according to a square function (Fig. 6a),
$\mu(t)$ $=\begin{cases}\mu_{m},&\text{if }|t-t_{0}|<\Delta_{t},\\[4.0pt] 0,&\text{otherwise},\end{cases}$
with the Gaussian pulse results obtained in the paper. The simulations reveal that both pulses achieve similar sigmoid-like probability transitions, indicating that it suffices to switch on the learning potential when a data sample is applied.
Figure 6: Single-site flip probability with different pulse shapes. (a) Numerical simulation of the transition probability when $\mu(t)$ varies according to a square pulse. (b) Numerical simulation of the switch probability when $\mu(t)$ follows a Gaussian pulse, reproduced from the main text for convenience.
@@ 2 $(0,0,1)$ $(0,0,1)$
.1 Simulation of learning under a spin-dependent local potential
Figure 7: Training of a 4x4 lattice with uncompensated interactions. (a, b) Complex transmissivity values for each of the $2^{16}$ possible spin configurations in the absence (a) and presence (b) of compensating terms. (c) Evolution of the real part of the transmissivity as the lattice is trained towards target transmissivities of 1-1 (blue), 0 (orange) and 11(green). (d) Final value of the real (blue) and imaginary (orange) parts of the transmissivity as the lattice is trained towards different real targets.
In Table 1, only two of the coupling terms are essential for learning. The other five are introduced to compensate for the spin-dependent shifts on the natural frequencies of the various modes, i.e., the term ensures that the frequency of the computational modes $\psi_{x/y}$ is independent of the spin configuration. In Figure 7, we examine the learning dynamics when such compensation is not in place, and the local potential $\omega_{l}$ is dependent on the spin texture. Under these conditions, the output signal can take an arbitrary complex phase (Fig. 7a), differing from the original scenario where the transmitted amplitude had a fixed phase relation with the lattice excitation (Fig. 7b). Simulating the learning dynamics as the system is trained towards a fixed reference transmissivity reveals that the system is still able to learn (Fig. 7c), although the number of iterations that it takes to converge to the correct transmittance increases by a significant amount. We attribute this increase to the fact that the system presents fewer spin configurations that result in the required transmissivity value—as transmissivity values are spread around the complex plane; therefore, it takes longer to settle in the correct configuration. Remarkably, this system can be trained to negative transmissivities (Fig. 7c, d), potentially relaxing the need for inverse inputs in the Iris dataset test.
These results motivate us to explore whether the system can be trained beyond signal inversion, to produce signals that are phase-shifted by 90 degrees. We observe that this is indeed the case (Fig. 8a, b). These results indicate that, although presenting slower learning dynamics, removing the compensating terms from Table 1 also increases expressivity.
Figure 8: Training a 4x4 lattice to produce phase-shifted signals. (a) Evolution of the imaginary part of the transmissivity as the lattice is trained towards targets of 0.75i-0.75i (blue), 0 (orange) and $0.75i$ (green). (b) Final value of the real (blue) and imaginary (orange) parts of the transmissivity as the lattice is trained towards different imaginary targets.
Selection of resonator eigenmodes
Figure 9: Mapping model degrees of freedom to resonator eigenmodes. a Possible mode selection for the weight ($\sigma$) and computational ($\psi_{x}$ and $\psi_{y}$) degrees of freedom. Note that for the computational modes, the assignment between variables and local modes (gauge) varies between even and odd sites. The black dot indicates a region where the computational modes have opposing sign, and overlap with the weight mode. There, the nonlinear learning potential $H_{L}$ would be applied. b By placing the nonlinear coupling elements at the zeros of the computational modes, we can realize the simplified coupling interactions.
Here, we will investigate how a self-learning metamaterial could be constructed, following the simplified model discussed in the previous section. To realize a metamaterial with the right interactions, we need to identify mode pairs with alternating zeros at the boundary. These zeros allow for the placement of the coupler beams that realize individual interaction terms. Remarkably, the eigenpair with angular mode number $\ell=2$ (Fig. 9a) of a circularly-symmetric resonator naturally lends itself to realizing the required interaction—as both modes in the pair expose regions of zero field at the boundary with all four neighbors, and the coupler beams can be placed (Fig. 9b). Because these properties originate in the symmetry characteristics of the mode, and thus do not rely on specific physics of the system (photonics, MEMS, microwave…), this mode selection provides a path for the realizaton of self-learning materials in diverse platforms.
In Fig. 9a we have also highlighted the location for the application of the learning potential $H_{l}$, which couples the difference between computational modes $\psi_{x/y}$ to the learning mode $\sigma$. By selecting a learning mode with radial mode number $n>0$, we also construct regions where only one mode is nonzero, as well as regions where only two of the modes is nonzero (for any pair of modes). This allows the design to incorporate tuning electrodes for specific modes, as well as couplers for specific pairs.
While these modes correspond to free-standing drums, in weakly-coupled resonators the mode shape can be assumed to remain constant up to first order of perturbation theory [62, 63], and differences can be compensated systematically by editing the resonator geometry.
Testing the hypothesis that oscillator lattices can learn
This article provides evidence for the hypothesis that an oscillator network can learn to solve a classification problem. Because the site flips are probabilistic, there is always a possibility that the observed improvement in classification accuracy is a random occurrence and does not represent genuine learning. To determine the strength of the results, we perform a paired t-test, which allows us to determine the p-value, the probability to randomly obtain (null hypothesis) a final distribution at least as extreme as the result from the simulations. The p-value is calculated from a t-value, given by
$t=\frac{<d>\sqrt{n}}{\sigma(d)},$ (15)
Where $d_{i}$ is the difference between initial and final classification accuracies, and $\sigma(d)$ represents its sample standard deviation. Evaluating Eq. 15 for the initial and final data from Fig. 4c results in a t-value of 180.01, with 12800 samples.
For very large number of samples, the t-distribution is approximately Gaussian. Hence, the corresponding p-value is given by
$p\approx\frac{1}{t\sqrt{(}2\pi)}e^{-t^{2}/2},$ (16)
These results indicate that the probability of randomly obtaining a performance improvement as extreme as the one from Fig. 4c is on the order of $e^{-16200}$. These results provide very strong evidence for the presence of learning: If the null hypothesis were true, the probability to obtain such an improvement would be approximately $10^{-7035}$.